The accumulated amount that you will receive after 38 months if you deposit R13 300 into an account where money is worth 11,35% per year, compounded every two months is


[1] R15 690,19.

[2] R18 080,24.

[3] R14 117,08.

[4] R18 865,83

.[5] R18 988,31.


1
Expert's answer
2021-05-21T03:42:45-0400

"A=P(1+\frac{r}{i})^{n}"

"=13300(1+\frac{11.36}{6})^{\frac{38}{12}\times 6}"

"=18988,3093"

"=R18,988,31"